Compute exact answers mentally by finding compatible numbers or substitutions.
Show your work and explain which method you used.
a. 237 + 140 +23
b. 4 x 273 x 25
c. 11 x 28

Answer by kevwill(135)

Here's how I would do each problem:

a) First, add 237 and 23 to get 260, then add 140 to get 400.

b) First, multiply 4 and 25 to get 100, then multiply by 273 to get 27,300

c) Apply the distributed property in reverse: 11 * 28 = (10 + 1) * 28 = 280 + 28 = 308